Shift report 20020108 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 08 2002 - 15:56:15 EST

  • Next message: Flemming Videbaek: "Fw: access wed Jan 09"

    ------------------------------------------------------------
    Date: Tue Jan  8 15:56:15 2002, Shift: 08:00-16:00
    Supervisor: KH, with: FV
    Log book#12 pages 26 - 28
    
    Shift Summary:
    ------------------------------------------------------------
    Machine Status:
    No real beam during shift.
    MCR informed that dump beam voting page has been re-instituted.  That is now the way to let your 
    opinion regarding beam dumping be known.
    
    Achievements:
    1. Mainframe BB3 was replaced.  This "fixed" the trigger
       counters that have been tripping every few minutes.
    2. TPC and CAMAC pedestal runs.
    3. Started implementing electronics to have scaler counters
       of RHIC sorted scaler overlaps to plan polarization
       measurements for next year's run.
    
    Problems:
    1. The home disk on opus filled up.  This caused run
       6353 not to have the end of run data written into the
       run DB.  It also caused the DAQ to quit functioning for
       all practical purposes.  It was verified, however, that
       this run did get transferred to HPSS.  
       FV restored the proper files so that the DAQ control 
       could function properly again.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Jan 08 2002 - 15:56:54 EST